OpenStudy (anonymous):

For an Arc length S, area of sector A, and a central angle Theta of a circle of radius R, find the indicated quantity: A=29.4m^2 R=109.9m find Theta in radian measure

OpenStudy (anonymous):

any ideas @wolf1728 ?

OpenStudy (wolf1728):

Circle area = PI * r² Circle area = PI * 109.9² Circle area = 37,944.19 Sector area = 29.4 Sector Area / Circle Area = 0.0007748222 Central Angle (degrees) = 0.0007748222 * 360 Central Angle (degrees) =0.2789360031 Central Angle (radians) = 0.0048683517 that's a REALLY small angle

OpenStudy (anonymous):

haha yeah that is really small.

OpenStudy (wolf1728):

recheck my figures but that is what it seems to be.

OpenStudy (anonymous):

i just did. i can't figure out what could be wrong about it, everything seemed to come out as you showed.

OpenStudy (anonymous):

would you like to help me with a story problem?